OpenStudy (anonymous):

What if you have all C's an 1 D what GPA is that

OpenStudy (anonymous):

it depends how many C's you have

OpenStudy (anonymous):

Does it count semester test to? Becuz those r the only thig that brought my grades down

OpenStudy (anonymous):

I believe everything counts

OpenStudy (anonymous):

here is the idea A = 4.0 or 90 or above B = 3.0 or 80-89 C = 2.0 or 70-79 D = 1.0 or 60-69

OpenStudy (anonymous):

Oh okay this is what I got last year an I started a new school an trying to get permit an need higher than a 2.0

OpenStudy (anonymous):

An I'm new to the school I'm at now so we haven't had or full semester yet

OpenStudy (anonymous):

so if you have a class that is 4 hrs and you make a C then its 2.0 x 4 = 8 if its a 3 hr course then 3 * 2.0 = 6 now you add all the points and divide by the number of hours you have and that should get you your GPA

OpenStudy (anonymous):

I had 7 classes but my 7th class was a sport

OpenStudy (anonymous):

I can tell that you are in high school

OpenStudy (anonymous):

It's 8am to all the way to 3:15 but I stayed longer cuz I played a sprort

OpenStudy (anonymous):

im not too sure if college GPA and high school GPA are the same

OpenStudy (anonymous):

Oh okay:(

OpenStudy (anonymous):

Do you think till be 2.0 prbhigher

OpenStudy (anonymous):

like I said it depends how many C's you have but if you had a B in there and a D im certain you have at least a 2.0

OpenStudy (anonymous):

I didn't try the last bit of school whenever it was about to be summer so that hurt me

OpenStudy (anonymous):

I think I had one good grade an it was an A or B

OpenStudy (anonymous):

then that's good enough I think :) just keep at it

OpenStudy (anonymous):

Okay thank you

OpenStudy (anonymous):

Thank you!:)

OpenStudy (anonymous):

np :D